Let Z be a standard normal random variable.

P(c<=Z<=-0.93)=0.1395

Carry your intermediate computations to at least four decimal places. Round your answer to two decimal places.
Answer #1

In this PCOSz5-0.93)=0:1395 Plzs-0.93) - P(ZSC) = 0.1395 01762 - P(ZSC) = 0.1395 P(ESC) = 0.1762 -0.1395 PIESC) = 0.0367 c=1–